# EvRWKV: A Continuous Interactive RWKV Framework for Effective Event-Guided Low-Light Image Enhancement

### v2

@@ V Conclusion
In this paper, we proposed EvRWKV, a novel and effective framework for LLIE that leverages continuous cross-modal interaction between event streams and low-light images through dual-domain processing. Our approach uniquely integrates the RWKV architecture with the EISFE module, enabling adaptive and robust fusion of asynchronous event data and standard images. The Cross-RWKV module effectively captures fine-grained temporal and spatial dependencies, while the EISFE module jointly performs noise suppression and structural restoration via frequency-domain filtering and spatial deformable convolutions. Extensive experiments on multiple challenging real-world low-light datasets demonstrate that EvRWKV significantly outperforms existing state-of-the-art methods in both quantitative metrics and visual quality, effectively suppressing noise, enhancing contrast, and restoring fine details even under severe degradation such as motion blur and extreme darkness. Our dual-domain fusion and adaptive gating mechanisms successfully balance complementary information from event and image modalities, resulting in superior enhancement robustness and generalization. In future work, we plan to explore further improvements in cross-modal fusion strategies and extend the application of the EvRWKV framework to other challenging imaging tasks such as real-time video enhancement and dynamic scene understanding in extreme lighting conditions.

### v3

@@ V Conclusion
In this paper, we proposed EvRWKV, an effective framework for LLIE that establishes continuous cross-modal interaction between event streams and low-light images. The Cross-RWKV Module, leveraging a recurrent structure, maintains feature consistency from low-level textures to high-level semantics, directly addressing the information loss and representational bottleneck issues of prior paradigms. This is complemented by the EISFE module, which performs robust dual-domain fusion by jointly handling frequency-domain noise and spatial-domain alignment Extensive experiments validate our approach. EvRWKV significantly outperforms state-of-the-art methods across both quantitative metrics and visual comparisons, effectively restoring fine details while suppressing severe artifacts. In future work, we identify two specific and promising directions. A primary avenue is the extension of EvRWKV to real-time video enhancement, where the recurrent structure and efficiency of our framework provide a strong foundation for maintaining temporal consistency. A second, more challenging direction is to improve robustness under extreme low-light conditions where event streams become exceedingly sparse, which may require novel strategies such as integrating generative priors to compensate for the data scarcity.
[Uncaptioned image] Wenjie Cai is currently pursuing the B.E. degree in Artificial Intelligence with the School of Artificial Intelligence, Anhui University, Hefei, China. He serves as a Research Intern at the Anhui Provincial International Joint Research Center for Advanced Technology in Medical Imaging. His research interests include image and video processing, computer vision and deep learning.
[Uncaptioned image] Qingguo Meng received his B.Eng. degree in computer science and technology from Henan Polytechnic University, Jiaozuo, China, in 2022. He is currently working on his Ph.D. at the School of Artificial Intelligence, Anhui University in Hefei, China. His research directions are object tracking, low-light image enhancement, medical imaging, and biometrics.
[Uncaptioned image] Zhenyu Wang is currently pursuing the B.E. degree in Artificial Intelligence with the School of Artificial Intelligence, Anhui University, Hefei, China. His research interests include image processing and computer vision.
[Uncaptioned image] Xingbo Dong (Member, IEEE) received the B.S. degree from Huazhong Agriculture University, Wuhan, China, in 2014, and the Ph.D. degree from the Faculty of Information Technology, Monash University, Melbourne, VIC, Australia, in 2021. He held a post-doctoral position with Yonsei University, Seoul, South Korea, in 2022. He is currently a Lecturer with Anhui University, Hefei, China. His research interests include biometrics, medical imaging, and image processing.
[Uncaptioned image] Zhe Jin (Member, IEEE) obtained a Ph.D. in Engineering from Universiti Tunku Abdul Rahman Malaysia (UTAR). He is a Professor at the School of Artificial Intelligence, Anhui University, China. His research interests include Biometrics, Pattern Recognition, Computer Vision, and Multimedia Security. He has published over 70 refereed journals and conference articles, including IEEE Trans. IFS, SMC-S, DSC, PR. He was awarded the Marie Skłodowska-Curie Research Exchange Fellowship. He visited the University of Salzburg, Austria, and the University of Sassari, Italy, respectively, as a visiting scholar under the EU Project IDENTITY 690907.